OpenStudy (anonymous):

The average (arithmetic mean) of the 11 numbers in a list is 14. If the average of 9 of the numbers in the list is 9, what is the average of the other 2 numbers?

OpenStudy (cwrw238):

sum of the 11 numbers = 11*14 = 154 9* 9 = 81 sum of last two numbers = 154-81 = 73 average of these 2 = 73/2 = 36.5

OpenStudy (anonymous):

thanku

OpenStudy (cwrw238):

welcome